# SpikeDet: Better Firing Patterns for Accurate and Energy-Efficient Object Detection with Spiking Neuron Networks

### v3

@@ IV-A Overview
The architecture of our proposed SpikeDet is shown in Fig. 3. As illustrated, the model adopts the backbone-neck-head architecture widely used in object detection frameworks [17, 4, 8, 38, 40, 42]. For the backbone, we propose MDSNet to extract multi-scale features with more stable firing patterns. For the neck, we design a Spiking Multi-direction Fusion Module (SMFM), which enables the model to enhance backbone features through multi-direction fusion. For the detection head, as it focuses on utilizing already formed features to generate detection results, establishing the stability of the firing pattern is not its critical function. Therefore, we directly adopt the existing SpikeYOLO Detection Head [20].
The spike-driven characteristics of SNNs make them naturally suitable for event data. Consequently, SNN-based detectors are now being explored on not only traditional static datasets but also event-based datasets. To handle both types of data effectively and fully exploit the temporal dynamics of SNNs, we encode the inputs as $4D$ tensors $[T,C,H,W]$, where $T$ represents time steps, $C$ denotes the number of channels, while $H$ and $W$ correspond to spatial dimensions. For static data, we employ the direct coding method [48], replicating the input $T$ times to form a time train that supports the temporal processing requirements of the model. For event data, event cameras can asynchronously collect millions of events per second. Each event is defined as $e_{k}=(t_{k},p_{k},x_{k},y_{k})$, where $x_{k}$ and $y_{k}$ are spatial coordinates, $t_{k}$ is the timestamp, and $p_{k}$ represents polarity. Therefore, to effectively reduce data volume while preserving sparse characteristics, we adopt the coding approach from [25, 20]. This approach first selects a specific time window, divides it into $T$ temporal bins, and integrates events within each bin into frame format, thereby forming the aforementioned 4D tensor representation. Finally, we use Rate Decoding [17] for model output decoding, followed by Non-Maximum Suppression (NMS) post-processing.
@@ IV-B1 Motivation
Object detection is a complex task that demands strong feature extraction capabilities from the backbone network. In SNNs, to address this requirement, extensive studies have been conducted on SNN-friendly residual learning approaches [25, 34, 33], as constructing deep networks through residual learning has been proven beneficial in ANNs. Among these efforts, MS-ResNet [25] is the most successful approach, featuring its membrane-based shortcut. Building upon this, subsequent backbones such as EMS-ResNet [15] and S-Backbone [20] have been proposed, enabling the development of advanced SNN-based object detectors. Taking EMS-ResNet as an example, this network can be described as follows,
@@
$\mathbf{y}_{I}^{l}=\psi\left(\psi(\mathbf{x}^{l})\right)+\mathbf{x}^{l},$ (9)
$\mathbf{y}_{D1}^{l}=\psi\left(\psi(\mathbf{x}^{l})\right)+\mathrm{Cat}[\mathrm{MPool}(\mathbf{x}^{l}),\psi(\mathrm{MPool}(\mathbf{x}^{l}))],$ (10)
$\mathbf{y}_{D2}^{l}=\psi\left(\psi(\mathbf{x}^{l})\right)+\psi(\mathrm{MPool}(\mathbf{x}^{l})).$ (11)
@@
Here, $\mathbf{y}$ represents output and $\mathbf{x}$ represents input. $\mathrm{tdBN}$, $\mathrm{Conv}$, $\mathrm{SN}$, $\mathrm{MPool}$, and $\mathrm{Cat}$ denote threshold-dependent Batch Normalization [33], convolution, I-LIF neurons, maxpool, and channel dimension concatenation, respectively. The superscript $l$ denotes the $l$-th residual block, while subscripts $D$ and $I$ indicate whether the residual block performs downsampling or not. Accordingly, Eq. 9 represents the MS-Block without downsampling, while Eqs. 10 and 11 represent the downsampling blocks EMS-Block1 and EMS-Block2, respectively.
When the model goes deeper, MS-Blocks are stacked repeatedly. For ease, let $\mathbf{y}_{R}^{l}$ denote the residual path output and $\mathbf{y}_{S}^{l}$ denote the shortcut path output. We assume the distributions of these two variables to be $\mathbf{y}_{R}^{l}\sim\mathcal{N}(0,{\sigma_{R}^{l}}^{2})$ and $\mathbf{y}_{S}^{L}\sim\mathcal{N}(0,{\sigma_{S}^{l}}^{2})$, where ${\sigma_{R}^{l}}^{2}$ and ${\sigma_{S}^{l}}^{2}$ represent their respective variances. Assuming the two variables are approximately independent [49], $\mathbf{y}_{I}^{l}\sim\mathcal{N}(0,{\sigma_{R}^{l}}^{2}+{\sigma_{S}^{l}}^{2})$. Since there is no computation on the shortcut path, for $k$ stacked MS-Blocks, we can recursively derive that $\mathbf{y}_{I}^{l}\sim\mathcal{N}(0,{\sigma_{S}^{1}}^{2}+\sum_{l=1}^{k}{\sigma_{R}^{l}}^{2})$. Therefore, the variance of $\mathbf{y}_{I}^{l}$ becomes increasingly large as the network deepens. This results in unstable membrane synaptic input distribution in subsequent neurons.
Proposition 2. For I-LIF neurons, if the input follows $\mathbf{x}^{t,n-1}\sim\mathcal{N}(0,{\sigma_{in}}^{2})$, then the membrane potential follows $\mathbf{u}^{t,n}\sim\mathcal{N}(0,{\sigma_{out}}^{2})$, where ${\sigma_{out}}^{2}$ is proportional to ${\sigma_{in}}^{2}$.
Proof The detailed proof is presented in Appendix A.
With Proposition 2, we can conclude that the aforementioned variance accumulation in synaptic input directly creates unstable membrane potential distributions. This instability, according to Eq. 2, increases the probability of neurons firing large integer-valued spikes, thereby exacerbating the local firing saturation problem, severely impacting object detection tasks as discussed in Section I. Fig. 4 confirms our analysis, visually demonstrating that unstable input leads to more active neuron firing patterns. Additionally, as shown in Eqs. 10 and 11, during downsampling, the model positions Maxpool operations before I-LIF neurons. Since Maxpool selects maximum values from local regions, this increases the input intensity received by I-LIF neurons, raising the risk of firing saturation.
@@
Figure 4: Firing rate distribution of I-LIF neurons for different presynaptic input $\mathbf{x}^{t,n}$ when $T=1$ and $D=4$. Increased variance leads to higher firing saturation probability.
@@ IV-B2 MDSNet Architecture
To address these issues, as shown in Fig. 3, we propose MDSNet, a novel SNN-based ResNet-like backbone network. On the shortcut path, we propose a Membrane-based Deformed Shortcut (MDS), which incorporates I-LIF-1$\times$1Conv-tdBN (1$\times$1-LCB). This structure converts input to spikes through the I-LIF layer and processes them through a 1$\times$1 convolution, which improves expressiveness capability while allowing batchnorm-scale-fusion with tdBN during inference [33]. Finally, the tdBN component provides this path with the ability to adjust the output distribution. On the residual path, we place MS-Blocks in it rather than on the main path. It provides a clean gradient propagation path for deep networks. Additionally, to avoid the aforementioned problems with stacked MS-Blocks, we place 1$\times$1-LCBs at both the input and output to control variance. Through improvements on both paths, we not only ensure the output distribution stability of each block in MDSNet but also preserve the advantage of efficient information propagation in deep networks offered by the membrane-based shortcut through structure enhancements.
Specifically, MDSNet consists of MDS-Block1 and MDS-Block2. MDS-Block1 is designed without downsampling, following the structure described above. MDS-Block2 serves as the downsampling block. Its residual path is identical to MDS-Block1, but in the shortcut path, we position Maxpool after the I-LIF layer to prevent excessive I-LIF input. Additionally, to better preserve balanced information during downsampling, we also employ fixed stride downsampling. After downsampling, the results from stride downsampling and Maxpool are weighted and added before convolution, followed by tdBN for distribution adjustment. To eliminate non-spike computations caused by pre-convolution addition during inference, we reparameterize this structure by splitting the shortcut path. One path combines fixed stride downsampling with 1$\times$1 convolution to form a 1$\times$1-LCB path, while the other integrates Maxpool after I-LIF of 1$\times$1-LCB.
@@ IV-C Analysis of Gradient Vanishing/Explosion for MDSNet
To demonstrate that the MDS does not affect the training and convergence, we use the Block Dynamical Isometry [50] to verify that MDSNet can effectively overcome the gradient vanishing or explosion problem.
@@
Defination 1. (Definition 3.1. in [50]) Consider a neural network that can be represented as a series of blocks as Eq. 12 and the j-th block’s Jacobian matrix is denoted as $\mathbf{J_{j}}$. If $\forall j$, $\phi(\mathbf{J_{j}}\mathbf{J_{j}}^{T})\approx 1$ and $\varphi(\mathbf{J_{j}}\mathbf{J_{j}}^{T})\approx 0$, the network achieves Block Dynamical Isometry and can avoid gradient vanishing or explosion.
Definition 2. (Definition 5.1. in [50]) Let $\mathbf{f(x)}$ be a transform whose Jacobian matrix is $\mathbf{J}$. $\mathbf{f}$ is called general linear transform when it satisfies:
$E\left[\frac{||\mathbf{f(x)}||_{2}^{2}}{len(\mathbf{f(x)})}\right]=\phi\left(\mathbf{J}\mathbf{J}^{T}\right)E\left[\frac{||\mathbf{x}||_{2}^{2}}{len(\mathbf{x})}\right].$ (13)
Theorem 1. (Theorem 4.1. in [50]) Given $\mathbf{J}:=\prod_{j=L}^{1}\mathbf{J_{j}}$, where $\{\mathbf{J_{j}}\in\mathbb{R}^{m_{j}\times m_{j-1}}\}$ is a series of independent random matrices. If $(\prod_{j=L}^{1}\mathbf{J_{j}})(\prod_{j=L}^{1}\mathbf{J_{j}})^{T}$ is at least the $1^{st}$ moment unitarily invariant, we have
@@
Since the information transmitted in the network can be considered as random variables [51], $E\left[\frac{||\mathbf{x}||_{2}^{2}}{\text{len}(\mathbf{x})}\right]$ is regarded as the second-order moment of the input, denoted by $\alpha_{2}$. Building on the derivations in [50, 15, 33] and Definitions 1 and 2, we can independently analyze each MDS-Block. Subsequently, by applying Theorem 1 to these individual analyses, we can derive the overall behavior of the MDSNet.
@@
Proof. The details can be found in Appendix A.
As elucidated in [15, 25], in the majority of cases, $\phi(\mathbf{JJ}^{T})\approx 1$ is sufficient to guarantee that the network avoids gradient vanishing or explosion. The condition $\varphi(\mathbf{J_{j}}\mathbf{J_{j}}^{T})\approx 0$ is imposed to preclude the occurrence of accidental situations. Therefore, Proposition 3 provides theoretical support for the ability of MDSNet to address the problem of gradient vanishing or explosion.
@@ IV-D SMFM
Research in feature fusion for SNNs is limited to one-way and two-way fusion [17, 20], where large-scale features and small-scale features are combined only a limited number of times. This restricts the transmission of information within the model. For SNNs, this insufficient fusion leads to local firing saturation problems in the center of target regions and creates excessively high firing responses in irrelevant areas. To address these issues, we propose the Spiking Multi-direction Fusion Module (SMFM). As shown in Fig. 5, after one-way and two-way fusion, compared to the original feature map, the firing responses in irrelevant areas significantly increase, such as in the person’s legs. Moreover, after two-way fusion, the center region of the tennis racket exhibits local firing saturation. These issues are partially mitigated after three-way fusion and are successfully resolved with four-way fusion, which effectively combines multi-scale information while preserving firing patterns.
Figure 6: The architecture of proposed SF-Block. The difference between 3$\times$3-LDCB and 3$\times$3-LCB lies in that the former replaces the 3$\times$3 convolution with a 3$\times$3 depthwise convolution.
@@
To achieve upsampling/downsampling, we use Nearest-Neighbor Interpolation (NNI) and 3$\times$3 convolution with stride 2 as core elements, and combine these components with LCB. For feature fusion, we multiply the feature maps by learnable constants and then add them together. These constants can be fused with the previous tdBN layer, and since the features are presynaptic inputs that are directly fed into neurons, this method is SNN-friendly. In this way, we achieve better feature alignment through adaptive weighting of different-scale features without extra parameters. Then, as shown in Fig. 6, we propose the Spiking Fusion Block (SF-Block) to enhance fusion by processing post-addition features. This block consists of two sub-blocks. The first has two paths, with one containing two stacked 3×3-LCBs and another containing a 3×3-LDCB (with internal depthwise convolution) followed by a 1×1-LCB, which enables feature extraction from different receptive fields. The second is an MS-Block for residual learning that further adjusts these features.

### v4

@@ IV-A Overview
The architecture of our proposed SpikeDet is shown in Fig. 3. As illustrated, the model adopts the backbone-neck-head architecture widely used in object detection frameworks [fan_sfod_2024, zhao_detrs_2024, chen_yoloms_2025, yolov8_ultralytics, yolo11_ultralytics]. For the backbone, we propose MDSNet to extract multi-scale features with more stable firing patterns. For the neck, we design a Spiking Multi-direction Fusion Module (SMFM), which enables the model to enhance backbone features through multi-direction fusion. For the detection head, which primarily generates detection results from formed features, establishing firing pattern stability is not critical. Therefore, we directly adopt the existing SpikeYOLO Detection Head [luo_integervalued_2025].
The spike-driven characteristics of SNNs make them naturally suitable for event data. Consequently, SNN-based detectors are now being explored on not only traditional static datasets but also event-based datasets. To handle both types of data effectively and fully exploit the temporal dynamics of SNNs, we encode the inputs as $4D$ tensors $[T,C,H,W]$, where $T$ represents time steps, $C$ denotes the number of channels, while $H$ and $W$ correspond to spatial dimensions. For static data, we employ the direct coding method [kim_rate_2022], replicating the input $T$ times to form a time train that supports the temporal processing requirements of the model. For event data, event cameras can asynchronously collect millions of events per second. Each event is defined as $e_{k}=(t_{k},p_{k},x_{k},y_{k})$, where $x_{k}$ and $y_{k}$ are spatial coordinates, $t_{k}$ is the timestamp, and $p_{k}$ represents polarity. Therefore, to effectively reduce data volume while preserving sparse characteristics, we adopt the coding approach from [hu_advancing_2024, luo_integervalued_2025]. This approach first selects a specific time window, divides it into $T$ temporal bins, and integrates events within each bin into frame format, thereby forming the aforementioned 4D tensor representation. Finally, we use Rate Decoding [fan_sfod_2024] for model output decoding, followed by Non-Maximum Suppression (NMS) post-processing.
@@ IV-B1 Motivation
Object detection is a complex task that demands strong feature extraction capabilities from the backbone network. In SNNs, to address this requirement, extensive studies have been conducted on SNN-friendly residual learning approaches [hu_advancing_2024, fang_deep_2021, zheng_going_2021], as constructing deep networks through residual learning has been proven beneficial in ANNs. Among these efforts, MS-ResNet [hu_advancing_2024] is the most successful approach, featuring its membrane-based shortcut. This design enables efficient information propagation in deep SNN-based models. Building upon this, subsequent backbones such as EMS-ResNet [su_deep_2023] and S-Backbone [luo_integervalued_2025] have been proposed, enabling the development of advanced SNN-based object detectors. Taking EMS-ResNet as an example, this network can be described as follows,
@@
$\mathbf{y}_{I}^{t,l}=\psi\left(\psi(\mathbf{x}^{t,l})\right)+\mathbf{x}^{t,l},$ (8)
$\mathbf{y}_{D1}^{t,l}=\psi\left(\psi(\mathbf{x}^{t,l})\right)+\mathrm{Cat}[\mathrm{MPool}(\mathbf{x}^{t,l}),\psi(\mathrm{MPool}(\mathbf{x}^{t,l}))],$ (9)
$\mathbf{y}_{D2}^{t,l}=\psi\left(\psi(\mathbf{x}^{t,l})\right)+\psi(\mathrm{MPool}(\mathbf{x}^{t,l})).$ (10)
@@
Here, $\mathbf{y}$ represents the output and $\mathbf{x}$ represents the input. $\mathrm{tdBN}$, $\mathrm{Conv}$, $\mathrm{SN}$, $\mathrm{MPool}$, and $\mathrm{Cat}$ denote threshold-dependent Batch Normalization [zheng_going_2021], convolution, I-LIF neurons, Maxpool, and channel dimension concatenation, respectively. $\psi$ denotes I-LIF-Conv-tdBN (LCB). The superscript $l$ denotes the $l$-th residual block, while subscripts $D$ and $I$ indicate whether the residual block performs downsampling or not. Accordingly, Eq. 8 represents the MS-Block without downsampling, while Eqs. 9 and 10 represent the downsampling blocks EMS-Block1 and EMS-Block2, respectively.
When the model goes deeper, MS-Blocks are stacked repeatedly. For ease, let $\mathbf{y}_{R}^{t,l}$ and $\mathbf{y}_{S}^{t,l}$ denote the residual and shortcut path outputs, with variances $\text{Var}[\mathbf{y}_{R}^{t,l}]$ and $\text{Var}[\mathbf{y}_{S}^{t,l}]$, respectively. For the output variance of the MS-Block $\text{Var}[\mathbf{y}_{I}^{t,l}]=\text{Var}[\mathbf{y}_{R}^{t,l}+\mathbf{y}_{S}^{t,l}]$, we have
$\text{Var}[\mathbf{y}_{I}^{t,l}]=\text{Var}[\mathbf{y}_{R}^{t,l}]+2\text{Cov}[\mathbf{y}_{R}^{t,l},\mathbf{y}_{S}^{t,l}]+\text{Var}[\mathbf{y}_{S}^{t,l}].$ (11)
To further analyze this, we derive the following proposition. This property holds under standard initialization schemes [he2015delving]. It continues to hold approximately during and after training due to weight decay and the stabilizing effect of tdBN.
Proposition 1. For $d$ stacked LCB layers ($d\geq 1$) with arbitrary kernel sizes, under the convolution zero-mean weight and zero bias initialization, the output at the $(n+d)$-th layer $\mathbf{y}^{t,n+d}$ is uncorrelated with the input at the $n$-th layer $\mathbf{x}^{t,n}$.
Proof The detailed proof is presented in Appendix A-A.
Based on Proposition 1, the residual path $\mathbf{y}_{R}^{t,l}=\psi(\psi(\mathbf{x}^{t,l}))$ is the output of 2 stacked LCB layers processing the input $\mathbf{x}^{t,l}$, while the shortcut path $\mathbf{y}_{S}^{t,l}=\mathbf{x}^{t,l}$. Therefore, we have $\text{Cov}[\mathbf{y}_{R}^{t,l},\mathbf{y}_{S}^{t,l}]=0$. Thus $\text{Var}[\mathbf{y}_{I}^{t,l}]=\text{Var}[\mathbf{y}_{R}^{t,l}]+\text{Var}[\mathbf{y}_{S}^{t,l}]$. For $k$ stacked MS-Blocks, we can recursively derive that
$\text{Var}[\mathbf{y}_{I}^{t,k}]=\text{Var}[\mathbf{y}_{S}^{t,1}]+\sum_{l=1}^{k}{\text{Var}[\mathbf{y}_{R}^{t,l}]}.$ (12)
Therefore, the variance of $\mathbf{y}_{I}^{t,l}$ becomes increasingly large as the network deepens. This results in unstable membrane synaptic input distribution in subsequent neurons.
@@
Figure 4: Firing rate distribution of I-LIF neurons for different presynaptic input $\mathbf{x}^{t,n}$ when $T=1$ and $D=4$. The input $\mathbf{x}$ is sampled from Gaussian distributions with different variances. Increased variance leads to higher firing saturation probability.
Proposition 2. For I-LIF neurons, the variance of the membrane potential is proportional to the variance of the input, i.e., $\text{Var}[\mathbf{u}^{t,n}]\propto\text{Var}[\mathbf{x}^{t,n}]$.
Proof The detailed proof is presented in Appendix A-B.
With Proposition 2, we can conclude that the aforementioned variance accumulation in synaptic input directly creates unstable membrane potential distributions. According to Eq. 2, this instability increases the probability of neurons firing large integer-valued spikes, thereby exacerbating the local firing saturation problem. This severely impacts object detection tasks as discussed in Section I. Fig. 4 confirms our analysis, visually demonstrating that unstable input leads to more active neuron firing patterns. Additionally, as shown in Eqs. 9 and 10, during downsampling, the model positions Maxpool operations before I-LIF neurons. Since Maxpool selects maximum values from local regions, it increases the input intensity received by I-LIF neurons. This raises the risk of local firing saturation.
@@ IV-B2 MDSNet Architecture
To address these issues, as shown in Fig. 3, we propose MDSNet, a novel SNN-based ResNet-like backbone network. On the shortcut path, we propose a Membrane-based Deformed Shortcut (MDS), which incorporates 1×1-LCB, i.e., LCB using $\times1$ convolution. This structure converts input to spikes through the I-LIF layer and processes them through a $\times1$ convolution, which improves expressiveness capability. The tdBN component further provides this path with the ability to adjust the output distribution. This fundamentally addresses the variance accumulation problem in Eq. 12, which is also experimentally verified in Section V-C1. On the residual path, we place MS-Blocks in it rather than on the main path. It provides a clean gradient propagation path for deep networks. Additionally, to avoid the aforementioned problems with stacked MS-Blocks, we place a $\times1$-LCB at both the input and output to control variance. Through improvements on both paths, we ensure the output distribution stability of each block in MDSNet. Meanwhile, structural enhancements preserve the efficient information propagation advantage offered by the membrane-based shortcut in deep networks.
Specifically, MDSNet consists of MDS-Block1 and MDS-Block2. MDS-Block1 is designed without downsampling, following the structure described above. MDS-Block2 serves as the downsampling block. Its residual path is identical to MDS-Block1. In the shortcut path, however, we position Maxpool after the I-LIF layer to avoid the amplified input to I-LIF. Additionally, to better preserve balanced information during downsampling, we also employ fixed stride downsampling. After downsampling, the stride and Maxpool results are weighted, added, convolved, and followed by tdBN for distribution adjustment. To eliminate non-spike computations caused by pre-convolution addition during inference, we reparameterize this structure by splitting the shortcut path. One path combines fixed stride downsampling with $\times1$ convolution to form a $\times1$-LCB path, while the other integrates Maxpool after the I-LIF layer in $\times1$-LCB.
@@
Figure 6: The architecture of the proposed SF-Block. The difference between $\times3$-LDCB and $\times3$-LCB lies in that the former replaces the $\times3$ convolution with a $\times3$ depthwise convolution.
@@ IV-C Analysis of Gradient Vanishing/Explosion for MDSNet
To demonstrate that the MDS does not affect the training and convergence, we use the Block Dynamical Isometry [chen_comprehensive_2020] to verify that MDSNet can effectively overcome the gradient vanishing or explosion problem.
@@
Definition 1. (Definition 3.1. in [chen_comprehensive_2020]) Consider a neural network that can be represented as a series of blocks as Eq. 13 and the j-th block’s Jacobian matrix is denoted as $\mathbf{J_{j}}$. If $\forall j$, $\phi(\mathbf{J_{j}}\mathbf{J_{j}}^{T})\approx 1$ and $\varphi(\mathbf{J_{j}}\mathbf{J_{j}}^{T})\approx 0$, the network achieves Block Dynamical Isometry and can avoid gradient vanishing or explosion.
Definition 2. (Definition 5.1. in [chen_comprehensive_2020]) Let $\mathbf{f(x)}$ be a transform whose Jacobian matrix is $\mathbf{J}$. $\mathbf{f}$ is called general linear transform when it satisfies:
[…]
Theorem 1. (Theorem 4.1. in [chen_comprehensive_2020]) Given $\mathbf{J}:=\prod_{j=L}^{1}\mathbf{J_{j}}$, where $\{\mathbf{J_{j}}\in\mathbb{R}^{m_{j}\times m_{j-1}}\}$ is a series of independent random matrices. If $(\prod_{j=L}^{1}\mathbf{J_{j}})(\prod_{j=L}^{1}\mathbf{J_{j}})^{T}$ is at least the $1^{st}$ moment unitarily invariant, we have
@@
[…]
@@
Proof. The details can be found in Appendix A-C.
As elucidated in [su_deep_2023, hu_advancing_2024], in the majority of cases, $\phi(\mathbf{JJ}^{T})\approx 1$ is sufficient to guarantee that the network avoids gradient vanishing or explosion. The condition $\varphi(\mathbf{J_{j}}\mathbf{J_{j}}^{T})\approx 0$ is imposed to preclude the occurrence of accidental situations. Therefore, Proposition 3 provides theoretical support for the ability of MDSNet to address the problem of gradient vanishing or explosion.
@@ IV-D SMFM
Research in feature fusion for SNNs is limited to one-way and two-way fusion [fan_sfod_2024, luo_integervalued_2025], where large-scale and small-scale features are combined only a limited number of times. As discussed in Section I, such insufficient fusion leads to local firing saturation in both object centers and regions with high local contrast. This makes multi-directional feature fusion more critical in SNNs than in ANNs. To address these issues, we propose the Spiking Multi-direction Fusion Module (SMFM). It enables the model to repeatedly integrate multi-scale features through iterative fusion, allowing more comprehensive feature fusion. Fig. 5 provides visualization evidence that multi-direction fusion effectively mitigates local firing saturation.
@@
To achieve upsampling/downsampling, we use Nearest-Neighbor Interpolation (NNI) with $\times1$-LCB and $\times3$-LCB with a stride of 2 as core elements. For feature fusion, we multiply the feature maps by learnable constants and then add them together. These constants can be fused with the previous tdBN layer. As the features serve as presynaptic inputs to neurons, this method is compatible with SNNs. In this way, we achieve better feature alignment through adaptive weighting of different-scale features without extra parameters. Then, as shown in Fig. 6, we propose the Spiking Fusion Block (SF-Block) to enhance the fused features. This block consists of two sub-blocks. The first has two paths, with one containing two stacked $\times3$-LCBs and another containing a $\times3$-LDCB (with internal depthwise convolution) followed by a $\times1$-LCB, which enables feature extraction from different receptive fields. The second is an MS-Block for residual learning that further adjusts these features.